jhonyy9 (jhonyy9):

- is this proof acceptably ? - If n is a natural number >= 3, prove that allways will be one natural naber ,,a" and a natural number ,,b" such that this statement n=a+b+1 is true - so than : assume there exists a natural number k such that k >= 3 and there exists a pair of natural numbers, a_k and b_k, such that (a_k + b_k + 1) = k. Let a_(k+1) = a_k. So, a_(k+1) is a natural number. Let b_(k+1) = (b_k) + 1. So, b_(k+1) is a natural number. (a_(k+1) + b_(k+1) + 1) = [a_k + ((b_k) + 1) + 1] = [(a_k + b_k + 1) + 1] = (k + 1). And (k + 1) is a natural number >

jhonyy9 (jhonyy9):

And (k + 1) is a natural number >= 3. So, (k + 1) is a natural number >= 3 such that there exists a pair of natural numbers, a_(k+1) and b_(k+1), such that (a_(k+1) + b_(k+1) + 1) = (k + 1) If there exists a natural number k such that k >= 3 and there exists a pair of natural numbers, a_k and b_k, such that (a_k + b_k + 1) = k, then, for every natural number n >= k, there exists a pair of natural numbers, a_n and b_n, such that a_n + b_n + 1 = n. Let a_3 = 1, which is a natural number. Let b_3 = 1, which is a natural number. (a_3 + b_3 + 1) = 3, which is a natural number >= 3. THUS, for every natural number n >= 3, there exists a pair of natural numbers, a_n and b_n, such that a_n + b_n + 1 = n.